OpenStudy (anonymous):

please help @iGreen @dan815 @EclipsedStar @rational @sleepyjess @TheSmartOne

OpenStudy (anonymous):

Sandy works at a clothing store. She makes $9 per hour plus earns 15% commission on her sales. She worked 80 hours over the last two weeks and had a total of $2,465 in sales before taxes. Which of the following is closest to how much she will earn in hourly wages and commission for those two weeks? $478 $720 $1,090 $370

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

please help

OpenStudy (michele_laino):

here I am

OpenStudy (anonymous):

can u please help me

OpenStudy (michele_laino):

after 80 hours she has earned 80*9=$720

OpenStudy (anonymous):

so it is b

OpenStudy (michele_laino):

we have to know the earning from commissions. We only know the percentage on each commisions on sales

OpenStudy (michele_laino):

we have to add this quantity: 2465*15/100= $369.75 So total earning is: 720+369.75=...?

OpenStudy (michele_laino):

what is: 720+ 369.75=...?

OpenStudy (michele_laino):

total earning is: $1089.75